OpenStudy (anonymous):

What is the solution to the square root of 6x+1|end root| +1=x?

OpenStudy (cwrw238):

sqrt(6x + 1 ) + 1 = x sqrt(6x+1) = x - 1 square both sides: 6x + 1 = ( x - 1)^2 now solve this quadratic equation

OpenStudy (cwrw238):

can u continue from here?

OpenStudy (anonymous):

I got 6x-1=x^2-2x+1, added 1 to both sides adn then 2x to both sides and got 8x=x^2 +2. I know I need to isolate X but once I do, do I square both sides or do I make the equation equal 0 and then factor?

OpenStudy (cwrw238):

make equation = 0

OpenStudy (cwrw238):

x^2 - 8x + 2 = 0

OpenStudy (anonymous):

Do I factor?

OpenStudy (cwrw238):

that will not factor. use the quadratic formula x = [-(-8) +/- sqrt (8^2 - 4* !* 2)] / 2

OpenStudy (cwrw238):

this will give you 2 solutions for x check fo an extraneous solution by substituting back in the original equation